tanA = 1-cosB/sinB then find the value of tan2A-tanB=?
a.0
b.1
c.1/2
d.1/4

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
TanA =2sin2 (B/2)/2sin(B/2)cos(B/2) =tan(B/2) now :- tan2A =2tan(A)/1-tan2(A) = 2tan(B/2)/1-tan2(B/2) = tan2 B/2 = tanB.
 
Hence
 
tan 2A – tanB = 0
 
Regards
Arun (askIITians forum expert)